# StylizedGS: Controllable Stylization for 3D Gaussian Splatting

### v1

@@ 3.3 Perceptual Control
With the previous components of our approach, we can achieve both rapid and detailed style transfer while preserving the original content. Here, we attempt to address the challenge users encounter in customizing the stylization process. Drawing inspiration from [22], we propose a series of strategies and loss functions for 3DGS stylization to control various perceptual factors, including color, scale, and spatial areas, aiming to enhance the customization capability of our approach.
Color Control. The color information within an image is a crucial perceptual aspect of its style. Unlike other stylistic elements such as brush strokes or dominant geometric shapes, color is largely independent and stands out as a distinct characteristic. As illustrated in Fig. 3, users may seek to preserve the original color scheme by only transferring the pattern style, or combine the pattern style of one style image with the color style of another. Consequently, our method provides users with independent control over color style during the stylization process. To achieve this, in our color transfer step, we can pre-color 3D scenes to align with any user-desired hues. Similar to the luminance-only algorithm proposed in [44], we then utilize the YIQ color space to separate the learning of luminance and color information. As shown in the definition of color style loss $\mathcal{L}_{style}^{color}$ in Eq. 10, we extract the luminance channel in the rendered view $\mathcal{I}_{render}^{Y}$ and the style image $\mathcal{I}_{style}^{Y}$ for the luminance-only style loss $L_{style}^{lum}$ calculation. Meanwhile, the RGB channels are retained for the content loss $\mathcal{L}_{content}$ calculation.
@@
Figure 4: Scale Control Results. Our method supports users to flexibly control the scale of the basic style elements, such as increasing or decreasing the density of Lego blocks in the example of last row.
Scale Control. The style scale, such as the thickness of brushstrokes or the density of the basic grains (as depicted in Fig. 4), is a foundational style element and plays a vital role in defining visual aesthetics. The stroke size, as a typical example of the style scale, is influenced by two factors: the receptive field of the convolution in the VGG network and the size of the style image input to the VGG network, as identified by [23]. Our exploration reveals that simply adjusting the receptive field is efficient for achieving scale-controllable stylization in 3DGS. We modulate the size of the receptive field by manipulating the selection of layers in the VGG network and adjusting the weights for each layer during the computation of the scale style loss $\mathcal{L}_{style}^{scale}$. As shown in Eq. 11, the style losses from different layer blocks are amalgamated to form $\mathcal{L}_{style}^{scale}$.
@@
Figure 5: Spatial Control Results. By specifying regions in a single view and the corresponding style images for different regions, users can transfer different styles to desired regions within a scene. The users can specify some points (in blue) to generate region masks (in black).
@@
Spatial Control. Traditional 3D style transfer methods typically apply a single style uniformly across the entire scene, resulting in the same stylization across the composition. However, there are scenarios where it becomes necessary to transfer distinct styles to different areas or objects. Users may also want to specify region-to-region constraints for their preferred style distribution within a scene, as illustrated in Fig. 5. Users only need to specify some points in the desired region of a single view and then our method utilizes SAM [45] to extract masks with those specified points to ease the interaction burden. Similarly, users can also specify certain areas in the style image to delineate style regions for partial style transfer. Additionally, we introduce a mask-tracking strategy that dynamically adjusts the position of input points based on area thresholds, ensuring the consistent generation of masks across various training views. More details can be found in the supplementary document.
@@
To provide spatial guidance during the stylization process, we introduce a spatial style loss, denoted as $\mathcal{L}_{style}^{spatial}$. Assuming the user specifies $r$ regions in a single view of a scene to be matched with $r$ style regions, our $\mathcal{L}_{style}^{spatial}$ is formulated as follows:
$\mathcal{L}_{style}^{spatial}=\sum_{r}w_{r}\cdot\mathcal{L}_{style}(F_{b_{s}}(M_{r}^{c}\circ\mathcal{I}_{render}),F_{b_{s}}(M_{r}^{s}\circ\mathcal{I}_{style}^{r}))$ (12)
@@
Here, $M_{r}^{c}$ and $M_{r}^{s}$ represent the binary spatial masks on the rendered views and the $r$-th style image, respectively. Additionally, during the color match step, we transform the color information in the region of the style image to the corresponding masked area of the content image to improve color correspondence.
@@
Figure 7: Qualitative comparisons with the baseline methods on LLFF and T&T datasets. It can be seen that our method has better results which faithfully capture both the color styles and pattern styles across different views.
Figure 8: Qualitative comparisons with baselines on LLFF dataset. Compared to other methods, our method excels in learning intricate and accurate geometric patterns while effectively preserving the semantic content of the original scene.
4 Experiments
Datasets. We conduct extensive experiments on multiple real-world scenes from LLFF [46] which contains forward-facing scenes and Tanks & Temples (T&T) dataset [47] which includes unbounded ${}^{\circ}$ large outdoor scenes. The scene data is also employed to train photo-realistic 3DGS representation before stylization with official configuration. Furthermore, our experiments involve a diverse set of style images from [13], allowing us to assess our method’s capability to handle a wide range of stylistic exemplars.
Baseline. We compare our method with the state-of-the-art 3D stylization methods, ARF [13], Stylescene [7], StyleRF [48] and Ref-NPR [42] based on different 3D representations. Specifically, Stylescene [7] adopts point clouds featured by VGG features averaged across views as a scene representation and transforms the point-wise features by modulating them with the encoding vector of a style image for stylization. ARF [13] is a NeRF-based approach that performs nearest neighbor feature matching on VGG feature maps along with deferred back propagation to memorize efficiently optimized parameters. StyleRF [48] implements style transfer within the feature space of radiance fields in a zero-shot manner. Ref-NPR [42] facilitates controllable scene stylization utilizing radiance fields to stylize a 3D scene, with a single stylized 2D view taken as reference. For all methods, we use their released code and pre-trained models. As Ref-NPR is a reference-based method, we use AdaIN [49] to obtain a stylized reference view according to its paper.
Implementation Details. We implement StylizedGS based on 3DGS representation [17]. In the color match step, we fix the density component of the initial 3DGS and optimize it using $\mathcal{L}_{rec}$ for 200 iterations. During the stylization process, we fine-tune all parameters of $G_{\theta}^{sty}$ over 800 iterations, while disabling adaptive density control of 3DGS. We adopt a pretrained VGG-16 network [28] to extract feature maps $F_{render}$, $F_{style}$, and $F_{content}$ used in Eq. 3. The network consists of 5 convolution blocks and we use the ‘conv2’ and ‘conv3’ blocks as the feature extractors empirically.
The loss weights $(\lambda_{rec},\lambda_{sty},\lambda_{con},\lambda_{sca},\lambda_{opa},\lambda_{tv})$ are respectively set to $(0.2,2,0.005,0.05,0.05,0.02)$. The depth preservation loss weight $\lambda_{dep}$ is set to $0.01$ for all forward-facing captures and $0.05$ for all ${}^{\circ}$ captures. At each stylization training iteration, we render an image from a sampled view different from all the training views to compute losses. The Adam optimizer [50] is adopted with a learning rate exponentially decayed from $11e-$ 1 to $21e-$ 2.
4.1 Qualitative Comparisons
We show visual comparisons with Stylescene and ARF methods in Fig. 7. We can see that our method exhibits a better style match to the style image compared to the others. For instance, Stylescene tends to generate over-smoothed results and lacks intricate structures in the Truck scene, while ARF struggles with propagating the style consistently across the entire scene and exhibits deficiencies in color correspondence. In contrast, our method faithfully captures both the color tones and brushstrokes of “The Scream” throughout the entire scene, demonstrating a superior style matching compared to the baselines.
Fig. 8 presents additional comparison results between ARF, StyleRF, Ref-NPR, and our method on LLFF datasets. By jointly optimizing the density and color components, our method excels in learning intricate geometric patterns and meso-structure. Specifically, in the orchids scene with a Lego style, other methods captures only the color style or a basic square pattern, whereas our approach successfully illustrates the detailed structure of Lego blocks, clearly showing even the small cylindrical patterns. This is also evident in the horns scene featuring a snow style, where our method produces more visually pleasant results that faithfully match the style of the hexagonal snowflake structure. Our method also preserves more semantic information in the stylized scene as we utilize the depth map to keep the spatial distribution in an image.
TABLE I: Quantitative comparisons on stylization under novel views. We report ArtFID, SSIM, average training time (Avg. train), and average rendering FPS (Avg. FPS) for our method and other baselines. ‘-’ denotes StyleRF does not require individual stylization training.
Metrics ArtFID($\downarrow$) SSIM($\uparrow$) Avg. train($\downarrow$) Avg. FPS($\uparrow$)
Stylescene 52.75 0.13 6.37 min 0.71
StyleRF 40.61 0.41 - 5.8
ARF 35.73 0.29 1.78 min 8.2
Ref-NPR 33.56 0.31 2.31 min 7.3
Ours 28.29 0.55 0.87 min 153
4.2 Quantitative Comparisons
Another significant advantage of the 3DGS representation is that it can produce high-quality renderings in real time and greatly reduce training time. Therefore, we conducted quantitative comparisons with other methods to demonstrate both the high fidelity and efficiency of our approach. As the optimization time during the stylization process is notably influenced by the resolution, we randomly selected 50 stylized scenes with consistent resolution from the LLFF dataset and evaluated the average times. For the stylization metrics, we choose ArtFID [51] to measure the quality of stylization following [52] and SSIM (Structural Similarity Index Measure) between the content and stylized images to measure the performance of detail preservation following [33]. As shown in Tab. I, our method excels in both stylized visual quality and content preservation. Regarding efficiency, our method has significant advantages in average rendering FPS (Avg. FPS) and can achieve real-time free-view synthesis. At the same time, it also outperforms other optimization-based stylization methods in terms of average training time (Avg. train) for stylization and can achieve per scene stylization within 1 minute on a single NVIDIA RTX 4090 GPU. Note that StyleRF is a zero-shot method and does not require individual stylization training for each case, but its rendering efficiency is worse and cannot be viewed in real time.
User Study We also perform a user study among our method and other comparison methods. We randomly chose 20 sets of stylized views of the 3D scenes from both the LLFF and T&T datasets processed by different methods and invited 25 participants (including 15 males and 10 females, aged from 18 to 50). Our questionnaire was thoughtfully designed to encompass diverse participants across various professions and age groups, minimizing the risk of echo chamber effects. Participants were granted ample time for unbiased evaluation, ensuring a comprehensive and fair assessment. First, we presented the participants with a style image, the original scene, as well as two stylized videos generated by our method and a random comparison method. Then we asked the participants to vote on the video based on four evaluation indicators: visual preference, style match level, content preservation level, and 3D consistency quality. We collected 1000 votes for all evaluating indicators and presented the result in Fig. 9 in the form of boxplot. Our method, StylizedGS, outperforms those of other methods across all evaluation indicators, resulting in visually preferable and consistent results. A p-value of 0.005 indicates the statistical significance of the obtained results.
4.3 Controllable Stylization Results
Our method enables users to control stylization in three ways, including color, scale, and spatial region. The corresponding controllable stylization results are shown in Figs. 3, 4 and 5, respectively. Fig. 6 further demonstrates our method’s sequential control capability, enabling precise and flexible adjustments across different conceptual factors. In addition to the direct stylization result, we apply different controls sequentially from left to right: preserving the color of the original scene, increasing the scale of the pattern style, and adopting spatial control to apply two styles. This flexibility allows for refined modifications of key parameters or stylistic attributes, tailored to specific aesthetic preferences or artistic requirements. This empowers users to achieve finely tailored adjustments, fostering enhanced artistic expression and creative exploration.
@@ 4.3 Controllable Stylization Results
TABLE II: Quantitative ablation study results. We report ArtFID and SSIM for our method and other ablation studies over 50 randomly chosen stylized cases.
@@ 4.4 Ablation Study
We perform ablation studies to validate our design choices. The original scene and input style image are presented in Fig. 10 (a). Without applying the recoloring procedure, the stylized scene in (b) shows inferior color correspondence with the style image. In Fig. 10 (d)(e), we illustrate the importance of our depth preservation loss in preserving the original scene’s geometry. Without applying depth loss, the background of the Horse scene disappears, and the Trex’s skeleton becomes blurrier. The heatmap in the top left corner depicts the difference between the rendered depth maps and ground truth depth maps. The significant disparities in the heatmap of (d) also underscore the effectiveness of our depth preservation loss. In Fig. 10 (c), we highlight the importance of fine-tuning both the color and density components in 3DGS, which is crucial for learning intricate style patterns. Neglecting the optimization of the density makes it challenging to capture stroke details in the sketch and shape patterns in the texture style. Therefore, a delicate equilibrium is established wherein fine-tuning the density component harmonizes with the depth preservation loss, thereby facilitating the transfer of intricate style details while safeguarding the integrity of the original geometry. The quantitative results of ablation studies are presented in Tab.  II which also validate the significant impact of each component on the performance of our proposed method.
Additionally, we explore the impact of the 3DGS filter, addressing potential floaters in the reconstructed scenes. The floaters become more noticeable after the style transfer process. To mitigate this, we introduce a pre-filtering process. Specifically, we filter out 3D Gaussians with opacity below a certain threshold before we perform the style transfer fine-tuning process. Without the pre-filtering process, the empty space will be given textures. Fig. 11 shows the comparisons and we highlight some noticeable local areas with red bounding boxes. The first row presents results without pre-filtering, where the stone’s and leaves’ textures extend into the air around the objects. In the second row, after pre-filtering and training, the scene exhibits enhanced clarity, with fewer visible colored floaters. Tab. II shows the quantitative results of the ablation studies which demonstrate that all components contribute to our proposed method.

### v2

@@ 3.3 Perceptual Control
Building on the previous components of our approach, we can achieve both rapid and intricate style transfer while preserving the original content. Here, we address the challenge users face in customizing the stylization process. Inspired by [24], we propose a series of strategies and loss functions for 3DGS stylization to control various perceptual factors, including color, scale, and spatial areas, enhancing the customization capabilities of our approach. The stylization loss will be replaced with a different form, namely $L_{style}^{color}$, $L_{style}^{scale}$ or $L_{style}^{spatial}$ defined below.
Color Control. Color information within an image is a crucial perceptual aspect of its style. Unlike other stylistic elements such as brush strokes or dominant geometric shapes, color is largely independent and stands out as a distinct characteristic. As illustrated in Fig. 3.2, users may seek to preserve the original color scheme by transferring only the pattern style or by combining the pattern style of one image with the color style of another. Consequently, our method provides users with independent control over color style during the stylization process. To achieve this, in our color transfer step, we can pre-color 3D scenes to align with any user-desired hues. Similar to the luminance-only algorithm proposed in [56], we utilize the YIQ color space to separate the learning of luminance and color information. As defined in the color style loss $\mathcal{L}_{style}^{color}$ in Eq. 9, we extract the luminance channel from the rendered view $\mathcal{I}_{render}^{Y}$ and the style image $\mathcal{I}_{style}^{Y}$ for the luminance-only style loss calculation. Meanwhile, the RGB channels are retained for the content loss $\mathcal{L}_{content}$ calculation.
@@
Figure 4: Scale Control Results. Our method enables users to flexibly control the scale of basic style elements, such as adjusting the density of Lego blocks, as demonstrated in the example provided in the last row.
Scale Control. The scale-related style features, such as the thickness of brushstrokes or the density of the basic grains (as depicted in Fig. 4), are foundational style elements and play a vital role in defining visual aesthetics. Here, the ‘scale’ we mention refers to the style pattern size. The stroke size, as a typical example of scale features, is influenced by two factors: the receptive field of the convolution in the VGG network and the size of the style image input to the VGG network, as identified by [25]. Our exploration reveals that simply adjusting the receptive field is efficient for achieving scale-controllable stylization in 3DGS. We modulate the size of the receptive field by manipulating the selection of layers in the VGG network and adjusting the weights for each layer during the computation of the scale style loss $\mathcal{L}_{style}^{scale}$. As shown in Eq. 10, the style losses from different layer blocks are combined as $\mathcal{L}_{style}^{scale}$:
@@
Figure 5: Spatial Control Results. By specifying masks in the content images, users can transfer different styles to desired regions. Users can input text prompt or specify certain points (highlighted in blue) to generate region masks (depicted in black).
@@
Spatial Control. Traditional 3D style transfer methods typically apply a single style uniformly across the entire scene, resulting in uniform stylization throughout the composition. However, there are scenarios where it becomes necessary to transfer distinct styles to different areas or objects. Users may also wish to specify region-to-region constraints for their preferred style distribution within a scene, as illustrated in Fig.5. We provide point-based interaction using SAM [57] and a proposed mask-tracking strategy, and a language-based approach using the LangSAM method [58] to generate masks across different views. Further details can be found in the supplementary document. Similarly, users can also specify certain areas in the style image to delineate style regions for partial style transfer.
@@
To provide spatial guidance during the stylization process, we introduce a spatial style loss, denoted as $\mathcal{L}_{style}^{spatial}$. Assuming the user specifies $r$ regions in a single view of a scene to be matched with $r$ style regions, the spatial loss $\mathcal{L}_{style}^{spatial}$ is formulated as follows:
$\mathcal{L}_{style}^{spatial}=\sum_{r}w_{r}\cdot\mathcal{L}_{style}(F(M_{r}^{c}\circ\mathcal{I}_{render}^{r}),F(M_{r}^{s}\circ\mathcal{I}_{style}^{r}))$ (11)
@@
Here, $M_{r}^{c}$ and $M_{r}^{s}$ represent the binary spatial masks on the rendered views and the $r$-th style image, respectively. Additionally, during the color transfer step, we transform the color information in the region of the style image to the corresponding masked area of the content image to improve color correspondence.
Figure 6: The User Interface of StylizedGS. Users can upload style and multi-view scene images via the right control panel, and the stylization results are displayed in the left panel in real time after optimization. The interface supports view control, color control, scale control, and spatial control with point-based and language-based interactions.
3.4 User Interface
We develop an interactive user interface integrated with our proposed method to facilitate convenient image-based scene stylization and various perceptual controls, as illustrated in Fig. 6. Users can upload a style image and multi-view scene images through the right control panel to generate stylized scenes. After the optimization, our interface displays the stylized scenes in the left window which can be viewed in real time. Additionally, the system supports interactive controls and various render modes by adjusting options in the control panel. Users can change the views by dragging the display window. For spatial control, we have implemented both point-based and language-based interactions to flexibly generate masks. Furthermore, users can customize the stylized scenes by adjusting the optimization steps and loss functions, combining different perceptual controls and style images, thereby supporting detailed and expressive stylization. We also provide a video of manipulating the interactive interface in the supplementary material.
4 Experiments
We include the implementation details, especially the training settings in our supplementary document.
Datasets. We conduct extensive experiments on multiple real-world scenes inLLFF [59] which contains forward-facing scenes and Tanks & Temples (T&T) dataset [60] which includes unbounded 360 large outdoor scenes. Furthermore, our experiments involve a diverse set of style images from [13] and WikiArt [61], allowing us to assess our method’s capability to handle a wide range of stylistic exemplars.
Baselines. We compare our method with the state-of-the-art 3D stylization methods, including LSNV [7], ARF [13], StyleRF [62], Ref-NPR [53] and StyleGaussian [22], which are based on different 3D representations. Specifically, LSNV is based on point cloud, ARF, StyleRF and Ref-NPR are based on NeRF, and StyleGaussian is based on 3DGS. LSNV, StyleRF and StyleGaussian are feed-forward-based, while others are optimization-based. For all methods, we use their released code and pre-trained models. As Ref-NPR is a reference-based method, we use AdaIN [54] to obtain a stylized reference view according to its paper.
Figure 7: Qualitative comparisons with baselines on LLFF dataset. Compared to other methods, our method excels in learning intricate and accurate geometric patterns while effectively preserving the semantic content of the original scene. (Please zoom in for better view.)
4.1 Qualitative Comparisons
We show visual comparisons on the LLFF dataset in Fig. 7. By jointly optimizing the opacity and color components, our method performs better in learning intricate geometric patterns and meso-structures. Specifically, for the scene of “trex” with a Lego style, other methods capture only the color style or a basic square pattern, whereas our approach successfully illustrates the detailed structure of the Lego blocks, clearly depicting even the small cylindrical patterns. This is also evident in the scene of ”horns” with a snow style, where our method produces visually more pleasing results that faithfully match the style of the hexagonal snowflake structure. Additionally, our method preserves more semantic information in the stylized scene by utilizing the depth map to maintain spatial distribution in the image. We show more stylization results on the LLFF dataset in Fig. 9.
Fig. 8 presents comparison results on the T&T dataset. It can be seen that our method achieves a better style match with the style image compared to the others. For instance, LSNV tends to produce overly smoothed results and lacks intricate structures, while ARF struggles to consistently transfer the style across the entire scene and exhibits deficiencies in color correspondence, such as the results of the second group. StyleGaussian does not faithfully transfer the style, resulting in blurry results. In contrast, our method accurately captures both the color tones and brushstrokes throughout the entire scene, demonstrating superior style matching compared to the baselines. We show more stylization results on the T&T dataset in Fig. 10.
Figure 8: Qualitative comparisons with the baseline methods on T&T datasets. It can be seen that our method has better results which faithfully capture both the color styles and pattern styles across different views. (Please zoom in for better view.)
TABLE I: Quantitative comparisons on stylization under novel views. We report ArtFID, SSIM, DISTS, average training time (Avg. train), and average rendering FPS (Avg. FPS) for our method and other baselines. ‘-’ denotes that the method does not require individual stylization training.
Metrics ArtFID($\downarrow$) SSIM($\uparrow$) DISTS($\downarrow$) Avg. train($\downarrow$) Avg. FPS($\uparrow$)
LSNV 52.75 0.13 0.33 - 0.71
StyleRF 40.61 0.41 0.30 - 5.8
ARF 35.73 0.29 0.31 2.25 min 8.2
Ref-NPR 33.56 0.31 0.33 2.54 min 7.3
StyleGaussian 37.92 0.35 0.30 - 142
Ours 28.29 0.55 0.26 0.87 min 153
Figure 9: Stylization results on LLFF dataset. Our method demonstrates robust performance, effectively adapting to various styles and diverse scenes.
Figure 10: Stylization results on T&T dataset. It can be seen that our method faithfully captures both the color styles and pattern styles across different views.
4.2 Quantitative Comparisons
A significant advantage of the 3DGS representation is the capability of rendering high-quality views in real-time and with much less training time. Therefore, we conducted quantitative comparisons with other methods to demonstrate both the high fidelity and efficiency of our approach. As the optimization time during the stylization process is notably influenced by the resolution, we randomly selected 50 stylized scenes with consistent resolution from the LLFF dataset and evaluated the average times. For the stylization metrics, we choose ArtFID [63] to measure the quality of stylization following [64]. We also use SSIM [65] and DISTS [66] between the content and stylized images to measure the performance of detail preservation and structural similarity following [35, 67, 68, 69, 70]. As shown in Tab. I, our method excels in both stylized visual quality and content preservation. Regarding efficiency, our method has significant advantages in average rendering FPS (Avg. FPS) and can achieve real-time free-view synthesis. At the same time, it outperforms other optimization-based stylization methods in terms of average training time (Avg. train) for stylization and can achieve per-scene stylization within 1 minute on a single NVIDIA RTX 4090 GPU. Note that while StyleRF and StyleGaussian are feed-forward methods and do not require individual stylization training for each case, it’s worth mentioning that StyleRF’s rendering is slow and cannot be viewed in real time.
User Study We conducted a user study comparing our method with other baseline methods. We randomly selected 20 sets of stylized views of 3D scenes from both the LLFF and T&T datasets, processed by various methods, and invited 25 participants. Participants were then asked to vote on the video based on four evaluation indicators: visual preference, style match level, content preservation level, and 3D consistency quality. More details are included in the supplementary document. We collected a total of 1000 votes for all evaluation indicators and presented the results in Fig. 11 in the form of boxplots. Our method, StylizedGS, outperformed other methods across all evaluation indicators, indicating visually preferable and consistent results. A p-value of 0.00001 demonstrates the statistical significance of the obtained results.
4.3 Controllable Stylization Results
Our method enables users to control stylization in three ways, including color, scale, and spatial region. The corresponding controllable stylization results are shown in Figs. 3.2, 4 and 5, respectively. Fig. 12 shows an example where a sequential series of controls are applied to a scene across different conceptual factors. In addition to the direct stylization result, we sequentially apply different controls from left to right: preserving the color of the original scene, increasing the scale of the pattern style, and adopting spatial control to apply two styles. Such flexibility allows for controllable refinement of key parameters or stylistic attributes, satisfying specific aesthetic preferences or artistic requirements and fostering enhanced artistic expression.
@@ 4.3 Controllable Stylization Results
TABLE II: Quantitative ablation study results. We report ArtFID and SSIM for our method and other ablations over 50 randomly chosen stylized cases.
@@ 4.4 Ablation Study
We conducted ablation studies to validate our design choices. The original scene and input style image are presented in Fig. 13 (a). When the recoloring procedure is not applied, as seen in (b), the stylized scene exhibits inferior color correspondence with the style image. In Fig. 13 (d)(e), we illustrate the importance of our depth preservation loss in preserving the original scene’s geometry. Without applying depth loss, the background of the Horse scene disappears, and the Trex’s skeleton becomes blurrier. The heatmap in the top left corner depicts the difference between the rendered depth maps and ground truth depth maps. The notable differences in the heatmap of (d) further emphasize the efficacy of our depth preservation loss. In Fig. 13 (c), we highlight the importance of fine-tuning both the color and density components in 3DGS, which is crucial for learning intricate style patterns. Neglecting the optimization of the density makes it challenging to capture stroke details and texture patterns in the style image. Therefore, a balance is established wherein fine-tuning the density component harmonizes with the depth preservation loss, thereby facilitating the transfer of intricate style details while keeping the integrity of the original geometry.
Additionally, we investigate how the proposed 3DGS filter affects the reconstructed scenes. Normally, floaters become more noticeable after the style transfer process. As shown in Fig. 14, the texture of stones and leaves extend into the air around the objects without the 3DGS filter. In contrast, the results of our method exhibit enhanced clarity, with fewer visible colored floaters. Tab. II shows the quantitative results of the ablation studies, demonstrating the contributions of all components of our method.